V) How would the contractionary monetary policy affect the domestic economy’s trade balance compa...

v) How would the contractionary monetary policy affect the domestic economy’s trade balance compared to trade deficits (Expansionary Fiscal Policy)

2) Consider the following representation of a model economy,

C = 4 + 0.7Y

P = 5W

I = 1 − 2i

W/Pe = 0.02Y

Q = 0

Ms = 600

X = 0

L(i, Y ) = P(5Y − 2i)

i) Find the Aggregate Supply equation for this economy.

ii) What is “medium run” full employment output in this economy?

iii) What is the “medium run” rate of interest in this economy?

iv) What is the price-level for this economy in this “medium run”?

v) If the price expectation in this economy is Pe = 25, what would the actual short-run price level be?
Answer #1

1) Domestic Trade balance in the country was immensely affected by the contractionary monetary policy. Let us discuss briefly the facts below.  

Contractionary monetary refers to the curtailing of money supply in to the economy and also injecting the slow nature of gradually reducing the generation of income. The Govt. usually increases the rate of interest. The high rate of interest leads to low investment. Thus in turn curtails the high level of consumption. The low investment towards production leads to less Foreign Exchange reserve. It results in the reducing value of the exchange rate. Atlast Trade balance results in More Imports and Less Exports. In contrary, Expansionary monetary nurishes the balance of trade by reducing the interest rate by paving the way for more income generation leads to more investment and the high increase in the exchange value.

2) i) Supply equation = M/P= 600/5W+I+Q

ii Median full run employment = C/I*P, where C refers to Consumer expenditure was divided into price fixed for all the goods.

iii) Median run rate of interest = I/W/Pe.

iv) Price level = Supply equation/Q*P

v) Short run price level will be Pe/Ms= 25/600= 0.41.